Courtney is 26 years older than Minili. In 10 years, Courtney will be twice as old as Minili. How old is each of them now?
I got C = 42 now
and M = 16 now
I set it up like this
2 (C-26) + 10 = C + 10
I was told by my teacher to set it up like this:
2 (C-26+10) = C+10
I cannot get an answer to show that she is 26 years older that way. Please let me know what I am doing wrong?

Answer by ankor@dixie-net.com(22740)

I think you got it mixed up, maybe this will help explain it:
Write an equation each statement
"Courtney is 26 years older than Minili."
C = M + 26
;
Their age in 10 years:
C+10
and
M+10
:
"In 10 years, Courtney will be twice as old as Minili."
C + 10 = 2(M+10
C + 10 = 2M + 20
C = 2M + 20 - 10
C = 2M + 10
:
How old is each of them now?
From the 1st statement, Substitute (M+26) for C, find M
M + 26 = 2M + 10
26 - 10 = 2M - M
16 = M: just like you said
:
C = M + 26
C = 16 + 26
C = 42, just like you said
;
Check our solution in the statement:
"In 10 years, Courtney will be twice as old as Minili.
52 = 2(26)
:
Not sure what your problem is, you had the right answer, hope this helps.